CSL^{A}TEX

Jaroslav Šnajdr  Zdeněk Wagner  Jiří Zlatuška  The L^{A}TEX3 Project

April 24, 1997

1 Hyphenation—file hyphen.cfg

1 ⟨∗hyphen⟩

 2 \ProvidesFile{hyphen.cfg}[1997/01/30 CsLaTeX]

\@@T1Codes All characters in hyphenation patterns must be letters, i.e. their \catcode must be 11. There must be also correct uc/lc codes set. First for the T1 encoding.

 3 \expandafter\def\csname @@T1Codes\endcsname{%
 4 \def\@tempa##1##2{%
 5  \@tempcnta##1\relax \@tempcntb##2\relax \@tempb
 6 }%
 7 \def\@tempb{%
 8  \ifnum\@tempcnta>\@tempcntb\else
 9  \@tempc\@tempcnta
10  \advance\@tempcnta\@ne
11  \expandafter\@tempb
12  \fi
13 }%
14 \def\@tempc##1{\catcode##1=11}%
15 \ifnum\inputlineno=\m@ne\else
16  \@tempa{"A0}{"BC}
17  \@tempa{"E0}{"FF}
18  \@tempa{"80}{"9C}
19  \@tempa{"C0}{"DF}
20 \fi
21 \def\@tempc##1{%
22  \count@=##1\advance\count@ by -"20
23  \uccode##1=\count@
24  \lccode##1=##1
25 }%
26 \@tempa{`\a}{`\z}
27 \ifnum\inputlineno=\m@ne\else
28  \@tempa{"A0}{"BC}
29  \@tempa{"E0}{"FF}
30 \fi
31 \def\@tempc##1{%
32  \count@=##1\advance\count@ by "20
33  \uccode##1=##1
34  \lccode##1=\count@
35  \sfcode##1=999
36 }%
37 \@tempa{`\A}{`\Z}
38 \ifnum\inputlineno=\m@ne\else
39  \@tempa{"80}{"9C}
40  \@tempa{"C0}{"DF}
41 \fi

42 \uccode`\^^Y=`\I     % dotless i
43 \lccode`\^^Y=`\^^Y   % dotless i
44 \uccode`\^^Z=`\J     % dotless j, ae in OT1
45 \lccode`\^^Z=`\^^Z   % dotless j, ae in OT1
46 \ifnum\inputlineno=\m@ne\else
47  \lccode`\^^9d=`\i    % dotted I
48  \uccode`\^^9d=`\^^9d % dotted I
49  \lccode`\^^9e=`\^^9e % d-bar
50  \uccode`\^^9e=`\^^d0 % d-bar
51 \fi
52 \let\@tempa=\@undefined
53 \let\@tempb=\@undefined
54 \let\@tempc=\@undefined

55 }

\@@IL2Codes The same for IL2 (ISO-8859-2).

56 \expandafter\def\csname @@IL2Codes\endcsname{%
57 \catcode225=11 \lccode225=225 \uccode225=193
58 \catcode193=11 \lccode193=225 \uccode193=193
59 \catcode228=11 \lccode228=228 \uccode228=196
60 \catcode196=11 \lccode196=228 \uccode196=196
61 \catcode232=11 \lccode232=232 \uccode232=200
62 \catcode200=11 \lccode200=232 \uccode200=200
63 \catcode239=11 \lccode239=239 \uccode239=207
64 \catcode207=11 \lccode207=239 \uccode207=207
65 \catcode233=11 \lccode233=233 \uccode233=201
66 \catcode201=11 \lccode201=233 \uccode201=201
67 \catcode236=11 \lccode236=236 \uccode236=204
68 \catcode204=11 \lccode204=236 \uccode204=204
69 \catcode237=11 \lccode237=237 \uccode237=205
70 \catcode205=11 \lccode205=237 \uccode205=205
71 \catcode229=11 \lccode229=229 \uccode229=197
72 \catcode197=11 \lccode197=229 \uccode197=197
73 \catcode181=11 \lccode181=181 \uccode181=165
74 \catcode165=11 \lccode165=181 \uccode165=165
75 \catcode242=11 \lccode242=242 \uccode242=210
76 \catcode210=11 \lccode210=242 \uccode210=210
77 \catcode243=11 \lccode243=243 \uccode243=211
78 \catcode211=11 \lccode211=243 \uccode211=211
79 \catcode244=11 \lccode244=244 \uccode244=212
80 \catcode212=11 \lccode212=244 \uccode212=212
81 \catcode246=11 \lccode246=246 \uccode246=214
82 \catcode214=11 \lccode214=246 \uccode214=214
83 \catcode224=11 \lccode224=224 \uccode224=192
84 \catcode192=11 \lccode192=224 \uccode192=192
85 \catcode248=11 \lccode248=248 \uccode248=216
86 \catcode216=11 \lccode216=248 \uccode216=216
87 \catcode185=11 \lccode185=185 \uccode185=169
88 \catcode169=11 \lccode169=185 \uccode169=169
89 \catcode187=11 \lccode187=187 \uccode187=171
90 \catcode171=11 \lccode171=187 \uccode171=171
91 \catcode250=11 \lccode250=250 \uccode250=218
92 \catcode218=11 \lccode218=250 \uccode218=218
93 \catcode249=11 \lccode249=249 \uccode249=217
94 \catcode217=11 \lccode217=249 \uccode217=217
95 \catcode252=11 \lccode252=252 \uccode252=220
96 \catcode220=11 \lccode220=252 \uccode220=220
97 \catcode253=11 \lccode253=253 \uccode253=221
98 \catcode221=11 \lccode221=253 \uccode221=221

 99 \catcode190=11 \lccode190=190 \uccode190=174
100 \catcode174=11 \lccode174=190 \uccode174=174

101 }

Macros for hyphenation switching when encoding or language is changed.

Token registers \@@front@@ and \@@end@@ contain commands, that will be later the contents of \@hyphenation macro (see its definition).

102 \newtoks\@@front@@
103 \newtoks\@@end@@
104 \bgroup
105 \count0=0

\DeclareLanguage Command \DeclareLanuage loads hyphenation patterns for respective languages and encodings. First control sequences that redefine \lang to \lang encoding are appended to \@@front@@ register. By command \setup@hyphenation are loaded patterns for all encoding noted in optional

 parameter of \DeclareLanguage.
106 \newcommand{\DeclareLanguage}[4][\undefined]{%
107 \def\c@ding{#1}\def\und@fin@d{\undefined}%
108 \ifx\c@ding\und@fin@d
109  \setup@hyphenation{#3}{\expandafter\@gobble\string#2}{#4}%
110 \else
111  \global\@@front@@=\expandafter{\the\@@front@@
112  \global\expandafter\expandafter\expandafter\def\expandafter
113  \expandafter\expandafter#2\expandafter\expandafter\expandafter
114  {\csname\expandafter\@gobble\string#2 \f@encoding\endcsname}}%
115  \def\one@arg##1,##2\end{%
116  \def\c@ding{##1}
117  \ifx\c@ding\und@fin@d\else
118  \setup@hyphenation[##1]{#3}{\expandafter\@gobble\string#2}{#4}%
119  \one@arg##2\end
120  \fi

121 }

122  \one@arg#1,\undefined,\end
123 \fi

124 }

\setup@hyphenation This command updates the encoding (by command \end@update), which causes control sequences like \'a to expand to right characters. After setting cat/uc/lc codes hyphenation patterns are loaded. Last we add commands to \@@end@@ for switching the \language register to right value when encoding is changed.

125 \newcommand{\setup@hyphenation}[4][\undefined]{%
126 \def\c@ding{#1}\def\und@fin@d{\undefined}%
127 \ifx\c@ding\und@fin@d
128  \def\c@ding{}
129 \else
130  \fontencoding{#1}\enc@update
131  \csname @@\c@ding Codes\endcsname
132 \fi
133 \enc@update
134 \InputIfFileExists{#2}%
135 {\message{^^JLoading #3 hyphenation patterns and exceptions}%
136  \global\expandafter\expandafter\expandafter
137  \def\expandafter\expandafter
138  \csname#3 \c@ding\endcsname\expandafter
139  {\expandafter\language\expandafter=\the\count0
140  #4\relax}%
141  \if\c@ding\relax\relax

142  \expandafter\global\expandafter
143  \let\csname#3\expandafter\endcsname\csname#3 \endcsname
144  \else\message{for #1}%
145  \global\expandafter\edef\csname#3 @\c@ding\endcsname
146  {\noexpand\language=\the\count0\noexpand\relax}%
147  \fi
148  \language=\count0}%
149 {\errhelp{The configuration for hyphenation is incorrectly
150  installed.^^J%
151  If you don't understand this error message you need
152  to seek^^Jexpert advice.}%
153 \errmessage{OOPS!! Hyphenation patterns file #2 for #3 not found!}}%
154 \if\c@ding\relax\relax\else
155 \global\@@end@@=
156 \expandafter\expandafter\expandafter
157  {\expandafter
158  \the\expandafter\@@end@@\expandafter
159  \ifnum\expandafter\language\expandafter=\the\count0\relax
160  \csname#3 @\f@encoding\endcsname
161  \fi}
162 \fi
163 \advance\count0 by 1\relax

164 }%

Here we define languages and encodings that are loaded into the format. By default we load english, czech and slovak patterns in the IL2 encoding.

165 ⟨hyphen⟩%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
166 ⟨hyphen⟩%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
167 \DeclareLanguage{\english}{hyphen.tex}{\lefthyphenmin=2
168  \righthyphenmin=3
169  \@splitrequestedfalse}%
170 \DeclareLanguage[IL2]{\czech}{czhyphen.tex}{\lefthyphenmin=2
171  \righthyphenmin=3
172  \@requesthyphens}%
173 \DeclareLanguage[IL2]{\slovak}{skhyphen.tex}{\lefthyphenmin=2
174  \righthyphenmin=3
175  \@requesthyphens}%
176 ⟨hyphen⟩%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
177 ⟨hyphen⟩%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
178 \egroup

\@hyphenation \@hyphenation macro is defined as contents of registers \@@front@@ and \@@end@@. This macro is executed at every change of encoding.

179 \@@front@@=\expandafter\expandafter\expandafter%
180  {\expandafter\the\expandafter\@@front@@\the\@@end@@}%
181 \expandafter
182  \def\expandafter
183  \@hyphenation\expandafter{%
184  \the\@@front@@}%
185 \@@front@@={}\@@end@@={}
186 \def\nolanguage{\lefthyphenmin=99\righthyphenmin=99\language=255
187  \@splitrequestedfalse\relax}
188 \newif\if@splithyphens
189 \newif\if@splitrequested

Macros for splitting hyphens

190 \def\minus{-}%
191 \bgroup\catcode`-\active\expandafter\egroup

192 \expandafter\def\noexpand-{%
193 \ifx\protect\relax
194 \ifhmode
195  \ifinner
196  \expandafter\expandafter\expandafter\expandafter
197  \expandafter\expandafter\expandafter-%
198  \else
199  \expandafter\expandafter\expandafter\expandafter
200  \expandafter\expandafter\expandafter\firstd@sh
201  \fi
202 \else%
203  \expandafter\expandafter\expandafter-%
204 \fi
205 \else
206 \expandafter\string\expandafter\string-%
207 \fi}%
208 \def\@nehyph@n{-}
209 \def\splithyph@n{\ifnum\expandafter\hyphenchar\the\font>-1
210  \discretionary{\char`-}{\char\hyphenchar\the\font}{\char`-}\else
211  \discretionary{\char`-}{\char`-}{\char`-}\fi}
212 \def\b@xhyph@n{\hbox{-}}
213 \def\tw@hyph@ns{--}
214 \def\thr@@hyph@ns{---}
215 \catcode `\- = \active
216 \def \firstd@sh    {\ifnextch@r -\then \sec@ndd@sh \else
217  \ifnum\expandafter\hyphenchar\the\font=`-
218  \ifnum\righthyphenmin>2
219  \if l\nextchar\b@xhyph@n  %   make sure we
220  \else                     % don't chop off `-li'
221  \splithyph@n
222  \fi
223  \else
224  \splithyph@n
225  \fi
226  \else
227  \@nehyph@n
228  \fi
229  \fii}
230 \def \sec@ndd@sh #1{\ifnextch@r -\then \thirdd@sh  \else \tw@hyph@ns\fii}
231 \def \thirdd@sh  #1{\ifnextch@r -\then \m@nyd@sh  \else
232  \thr@@hyph@ns\fii}
233 \def \m@nyd@sh   #1{\thr@@hyph@ns\firstd@sh}
234 \def \ifnextch@r #1\then #2\else #3\fii
235  {\def \c@mp@rech@rs
236  {\ifx #1\nextchar
237  \def \next {#2}%
238  \else
239  \def \next {#3}%
240  \fi
241  \next

242 }%

243  \futurelet \nextchar \c@mp@rech@rs

244 }

245 \catcode`-=12
246 \def\standardhyphens{\catcode`-=12 \@splithyphensfalse}%
247 \def\splithyphens{\if@splitrequested\catcode`-\active\fi
248  \@splithyphenstrue}%
249 \def\@requesthyphens{\if@splithyphens\catcode`-=\active\fi

250  \@splitrequestedtrue}
251 \let\@looseness\looseness
252 \def\looseness{%
253  \if@splithyphens
254  \standardhyphens
255  \afterassignment\splithyphens
256  \fi
257  \@looseness

258 }

259 \expandafter\fontencoding\expandafter U\expandafter
260  \fontencoding\expandafter{\f@encoding}\enc@update
261 \@hyphenation
262 \english

263 ⟨/hyphen⟩

 2 Fonts—file fonttext.cfg

This is copy of original fonttext.ltx with definitions for IL2 added.

264 ⟨∗fonttext⟩

265 \ProvidesFile{fonttext.ltx}[1997/01/30 CsLaTeX]
266 \let\@hyphenation=\relax

Macro \DeclareFontEncoding@ is changed to execute \@hyphenation at every encoding change. This can be also achieved by changing file xxxenc.def, but standard encoding definition file can’t be changed because of L^{A}TEX copyright.

267 \def\DeclareFontEncoding@#1#2#3{%
268 \expandafter
269 \ifx\csname T@#1\endcsname\relax
270  \def\cdp@elt{\noexpand\cdp@elt}%
271  \xdef\cdp@list{\cdp@list\cdp@elt{#1}%
272  {\default@family}{\default@series}%
273  {\default@shape}}%
274  \expandafter\let\csname#1-cmd\endcsname\@changed@cmd
275 \else
276  \@font@info{Redeclaring font encoding #1}%
277 \fi
278 \global\@namedef{T@#1}{#2\@hyphenation}%
279 \global\@namedef{M@#1}{\default@M#3}%

280 }

281 \input {omlenc.def}
282 \input  {t1enc.def}
283 \input {ot1enc.def}       % <- should come after T1 for speed

Definition of IL2 encoding.

284 \input {il2enc.def}
285 \input {omsenc.def}
286 \fontencoding{OT1}
287 \DeclareFontEncodingDefaults{}{}
288 \DeclareFontSubstitution{T1}{cmr}{m}{n}
289 \DeclareFontSubstitution{OT1}{cmr}{m}{n}
290 \begingroup
291 \nfss@catcodes
292 \input  {t1cmr.fd}
293 \input  {ot1cmr.fd}

IL2 once again.

294 \input  {il2cmr.fd}
295 \endgroup

296 \begingroup
297 \nfss@catcodes
298 \input {ot1cmss.fd}
299 \input {ot1cmtt.fd}
300 \endgroup
301 \DeclareErrorFont{OT1}{cmr}{m}{n}{10}
302 \newcommand\rmdefault{cmr}
303 \newcommand\sfdefault{cmss}
304 \newcommand\ttdefault{cmtt}
305 \newcommand\bfdefault{bx}
306 \newcommand\mddefault{m}
307 \newcommand\itdefault{it}
308 \newcommand\sldefault{sl}
309 \newcommand\scdefault{sc}
310 \newcommand\updefault{n}
311 \newcommand\encodingdefault{OT1}
312 \newcommand\familydefault{\rmdefault}
313 \newcommand\seriesdefault{\mddefault}
314 \newcommand\shapedefault{\updefault}

315 ⟨/fonttext⟩

3 Definition of the IL2 encoding—file il2enc.def

316 ⟨∗il2enc⟩

317 \ProvidesFile{il2enc.def}[1997/04/21 CSLaTeX]
318 \DeclareFontEncoding{IL2}{\uccode152=152 \lccode152=184
319  \uccode184=152 \lccode184=184
320  \uccode165=165 \lccode165=181
321  \uccode181=165 \lccode181=181
322  \uccode169=169 \lccode169=185
323  \uccode185=169 \lccode185=185
324  \uccode171=171 \lccode171=187
325  \uccode187=171 \lccode187=187
326  \uccode174=174 \lccode174=190
327  \uccode190=174 \lccode190=190}{}
328 \DeclareTextAccent{\"}{IL2}{127}
329 \DeclareTextAccent{\'}{IL2}{19}
330 \DeclareTextAccent{\.}{IL2}{95}
331 \DeclareTextAccent{\=}{IL2}{22}
332 \DeclareTextAccent{\^}{IL2}{94}
333 \DeclareTextAccent{\`}{IL2}{18}
334 \DeclareTextAccent{\~}{IL2}{126}
335 \DeclareTextAccent{\H}{IL2}{125}
336 \DeclareTextAccent{\u}{IL2}{21}
337 \DeclareTextAccent{\v}{IL2}{20}
338 \DeclareTextAccent{\r}{IL2}{23}
339 \DeclareTextCommand{\b}{IL2}[1]
340 {\oalign{\null#1\crcr\hidewidth\sh@ft{29}%
341  \vbox to.2ex{\hbox{\char22}\vss}\hidewidth}}
342 \DeclareTextCommand{\c}{IL2}[1]
343 {\setbox\z@\hbox{#1}\ifdim\ht\z@=1ex\accent24 #1%
344  \else{\ooalign{\unhbox\z@\crcr\hidewidth\char24\hidewidth}}\fi}
345 \DeclareTextCommand{\d}{IL2}[1]
346 {\oalign{\null#1\crcr\hidewidth\sh@ft{08}.\hidewidth}}
347 \DeclareTextSymbol{\AE}{IL2}{29}
348 \DeclareTextSymbol{\OE}{IL2}{30}
349 \DeclareTextSymbol{\O}{IL2}{31}

350 \DeclareTextSymbol{\ae}{IL2}{26}
351 \DeclareTextSymbol{\i}{IL2}{16}
352 \DeclareTextSymbol{\j}{IL2}{17}
353 \DeclareTextSymbol{\oe}{IL2}{27}
354 \DeclareTextSymbol{\o}{IL2}{28}
355 \DeclareTextSymbol{\ss}{IL2}{25}
356 \DeclareTextSymbol{\textemdash}{IL2}{124}
357 \DeclareTextSymbol{\textendash}{IL2}{123}
358 \DeclareTextSymbol{\textexclamdown}{IL2}{60}
359 \DeclareTextSymbol{\textquestiondown}{IL2}{62}
360 \DeclareTextSymbol{\textquotedblleft}{IL2}{92}
361 \DeclareTextSymbol{\textquotedblright}{IL2}{`\"}
362 \DeclareTextSymbol{\textquoteleft}{IL2}{`\`}
363 \DeclareTextSymbol{\textquoteright}{IL2}{`\'}
364 \DeclareTextCommand{\L}{IL2}
365 {\leavevmode\setbox0\hbox{L}\hbox to\wd0{\hss\char32L}}
366 \DeclareTextCommand{\l}{IL2}{{\char32l}}
367 \DeclareTextCompositeCommand{\r}{IL2}{A}
368 {\leavevmode\setbox0\hbox{h}\dimen@\ht0\advance\dimen@-1ex%
369  \rlap{\raise.67\dimen@\hbox{\char'27}}A}
370 \DeclareTextCommand{\SS}{IL2}{SS}
371 \DeclareTextCommand{\textdollar}{IL2}{{%
372  \ifdim \fontdimen\@ne\font >\z@
373  \slshape
374  \else
375  \upshape
376  \fi
377  \char`\$}}
378 \DeclareTextCommand{\textsterling}{IL2}{{%
379  \ifdim \fontdimen\@ne\font >\z@
380  \itshape
381  \else
382  \fontshape{ui}\selectfont
383  \fi
384  \char`\$}}
385 \DeclareTextComposite{\'}{IL2}{l}{'345}%
386 \DeclareTextComposite{\'}{IL2}{r}{'340}%
387 \DeclareTextComposite{\'}{IL2}{a}{'341}%
388 \DeclareTextComposite{\'}{IL2}{e}{'351}%
389 \DeclareTextComposite{\'}{IL2}{\i}{'355}%
390 \DeclareTextComposite{\'}{IL2}{i}{'355}%
391 \DeclareTextComposite{\'}{IL2}{o}{'363}%
392 \DeclareTextComposite{\'}{IL2}{u}{'372}%
393 \DeclareTextComposite{\'}{IL2}{y}{'375}%
394 \DeclareTextComposite{\'}{IL2}{L}{'305}%
395 \DeclareTextComposite{\'}{IL2}{R}{'300}%
396 \DeclareTextComposite{\'}{IL2}{A}{'301}%
397 \DeclareTextComposite{\'}{IL2}{E}{'311}%
398 \DeclareTextComposite{\'}{IL2}{I}{'315}%
399 \DeclareTextComposite{\'}{IL2}{O}{'323}%
400 \DeclareTextComposite{\'}{IL2}{U}{'332}%
401 \DeclareTextComposite{\'}{IL2}{Y}{'335}%
402 \DeclareTextComposite{\`}{IL2}{a}{'270}%
403 \DeclareTextComposite{\`}{IL2}{A}{'230}%
404 \DeclareTextComposite{\v}{IL2}{c}{'350}%
405 \DeclareTextComposite{\v}{IL2}{d}{'357}%
406 \DeclareTextComposite{\v}{IL2}{e}{'354}%
407 \DeclareTextComposite{\v}{IL2}{n}{'362}%

408 \DeclareTextComposite{\v}{IL2}{r}{'370}%
409 \DeclareTextComposite{\v}{IL2}{s}{'271}%
410 \DeclareTextComposite{\v}{IL2}{z}{'276}%
411 \DeclareTextComposite{\v}{IL2}{t}{'273}%
412 \DeclareTextComposite{\v}{IL2}{l}{'265}%
413 \DeclareTextComposite{\v}{IL2}{u}{'371}%
414 \DeclareTextComposite{\v}{IL2}{C}{'310}%
415 \DeclareTextComposite{\v}{IL2}{D}{'317}%
416 \DeclareTextComposite{\v}{IL2}{E}{'314}%
417 \DeclareTextComposite{\v}{IL2}{N}{'322}%
418 \DeclareTextComposite{\v}{IL2}{R}{'330}%
419 \DeclareTextComposite{\v}{IL2}{S}{'251}%
420 \DeclareTextComposite{\v}{IL2}{T}{'253}%
421 \DeclareTextComposite{\v}{IL2}{Z}{'256}%
422 \DeclareTextComposite{\v}{IL2}{L}{'245}%
423 \DeclareTextComposite{\v}{IL2}{U}{'331}%
424 \DeclareTextComposite{\^}{IL2}{o}{'364}%
425 \DeclareTextComposite{\^}{IL2}{O}{'324}%
426 \DeclareTextComposite{\"}{IL2}{a}{'344}%
427 \DeclareTextComposite{\"}{IL2}{o}{'366}%
428 \DeclareTextComposite{\"}{IL2}{u}{'374}%
429 \DeclareTextComposite{\"}{IL2}{A}{'304}%
430 \DeclareTextComposite{\"}{IL2}{O}{'326}%
431 \DeclareTextComposite{\"}{IL2}{U}{'334}%
432 \DeclareTextComposite{\r}{IL2}{u}{'371}%
433 \DeclareTextComposite{\r}{IL2}{U}{'331}%
434 \DeclareTextSymbol{\clqq}{IL2}{254}%
435 \DeclareTextCommand{\crqq}{IL2}%
436  {{\edef\@SF{\spacefactor\the\spacefactor}\char255 \@SF\relax}}%

437 ⟨/il2enc⟩

4 .fd files for CS-fonts

Created by changing .fd files ot1*.fd

438 ⟨∗il2cmr⟩

439 \ProvidesFile{il2cmr.fd}[1997/02/06 CsLaTeX font definitions]
440 \DeclareFontFamily{IL2}{cmr}{\hyphenchar\font45 }
441 \DeclareFontShape{IL2}{cmr}{m}{n}
442 {  <5> <6> <7> <8> <9> <10> <12> gen * csr
443  <10.95> csr10
444  <14.4>  csr12
445  <17.28><20.74><24.88>csr17}{}
446 \DeclareFontShape{IL2}{cmr}{m}{sl}

447 {

448  <5><6><7>cssl8%
449  <8> <9> gen * cssl
450  <10><10.95>cssl10%
451  <12><14.4><17.28><20.74><24.88>cssl12%
452  }{}
453 \DeclareFontShape{IL2}{cmr}{m}{it}

454 {

455  <5><6><7>csti7%
456  <8>csti8%
457  <9>csti9%
458  <10><10.95>csti10%
459  <12><14.4><17.28><20.74><24.88>csti12%
460  }{}
461 \DeclareFontShape{IL2}{cmr}{m}{sc}

463  <5><6><7><8><9><10><10.95><12>
464  <14.4><17.28><20.74><24.88> cscsc10
465  }{}
466 \DeclareFontShape{IL2}{cmr}{m}{ui}

467 {

468  <5><6><7><8><9><10><10.95><12>%
469  <14.4><17.28><20.74><24.88>csu10%
470  }{}
471 \DeclareFontShape{IL2}{cmr}{b}{n}

472 {

473  <5><6><7><8><9><10><10.95><12>
474  <14.4><17.28><20.74><24.88> csb10
475  }{}
476 \DeclareFontShape{IL2}{cmr}{bx}{n}

477 {

478  <5> <6> <7> <8> <9> gen * csbx
479  <10><10.95> csbx10
480  <12><14.4><17.28><20.74><24.88>csbx12
481  }{}
482 \DeclareFontShape{IL2}{cmr}{bx}{sl}

483 {

484  <5> <6> <7> <8> <9>
485  <10> <10.95> <12> <14.4> <17.28> <20.74> <24.88> csbxsl10
486  }{}
487 \DeclareFontShape{IL2}{cmr}{bx}{it}

488 {

489  <5> <6>  <7>  <8> <9>
490  <10> <10.95> <12> <14.4> <17.28> <20.74> <24.88> csbxti10
491  }{}
492 \DeclareFontShape{IL2}{cmr}{bx}{ui}
493  {<->ssub * cmr/m/ui}{}

494 ⟨/il2cmr⟩ 495 ⟨∗il2cmdh⟩

496 \ProvidesFile{il2cmdh.fd}[1997/02/06 CsLaTeX font definitions]
497 \DeclareFontFamily{IL2}{cmdh}{\hyphenchar\font45 }
498 \DeclareFontShape{IL2}{cmdh}{m}{n}{
499  <10> csdunh10
500  }{}

501 ⟨/il2cmdh⟩ 502 ⟨∗il2cmfib⟩

503 \ProvidesFile{il2cmfib.fd}[1997/02/13 CsLaTeX font definitions]
504 \DeclareFontFamily{IL2}{cmfib}{\hyphenchar\font45 }
505 \DeclareFontShape{IL2}{cmfib}{m}{n}{
506  <8> csfib8
507  }{}

508 ⟨/il2cmfib⟩ 509 ⟨∗il2cmfr⟩

510 \ProvidesFile{il2cmfr.fd}[1997/02/13 CsLaTeX font definitions]
511 \DeclareFontFamily{IL2}{cmfr}{\hyphenchar\font45 }
512 \DeclareFontShape{IL2}{cmfr}{m}{n}{%
513  <10> csff10
514  }{}
515 \DeclareFontShape{IL2}{cmfr}{m}{it}{%
516  <10> csfi10
517  }{}

518 ⟨/il2cmfr⟩ 519 ⟨∗il2cmss⟩

520 \ProvidesFile{il2cmss.fd}[1997/02/13 CsLaTeX font definitions]
521 \DeclareFontFamily{IL2}{cmss}{\hyphenchar\font45 }
522 \DeclareFontShape{IL2}{cmss}{m}{n}

523 {

524  <5><6><7><8>csss8
525  <9>csss9
526  <10><10.95>csss10
527  <12><14.4>csss12
528  <17.28><20.74><24.88>csss17
529  }{}
530 \DeclareFontShape{IL2}{cmss}{m}{it}
531  {<->sub*cmss/m/sl}{}
532 \DeclareFontShape{IL2}{cmss}{m}{sl}

533 {

534  <5><6><7><8>csssi8 <9> csssi9
535  <10><10.95>csssi10
536  <12><14.4>csssi12
537  <17.28><20.74><24.88>csssi17
538  }{}
539 \DeclareFontShape{IL2}{cmss}{m}{sc}
540  {<->sub*cmr/m/sc}{}
541 \DeclareFontShape{IL2}{cmss}{m}{ui}
542  {<->sub*cmr/m/ui}{}
543 \DeclareFontShape{IL2}{cmss}{sbc}{n}

544 {

545  <5> <6> <7> <8> <9> csssdc10
546  <10> <10.95> <12> <14.4> <17.28> <20.74> <24.88>csssdc10
547  }{}
548 \DeclareFontShape{IL2}{cmss}{bx}{n}

549 {

550  <5> <6> <7> <8> <9> csssbx10
551  <10> <10.95> <12> <14.4> <17.28> <20.74> <24.88> csssbx10
552  }{}
553 \DeclareFontShape{IL2}{cmss}{bx}{ui}
554  {<->sub*cmr/bx/ui}{}

555 ⟨/il2cmss⟩ 556 ⟨∗il2cmtt⟩

557 \ProvidesFile{il2cmtt.fd}[1997/02/13 CsLaTeX font definitions]
558 \DeclareFontFamily{IL2}{cmtt}{\hyphenchar \font\m@ne}
559 \DeclareFontShape{IL2}{cmtt}{m}{n}

560 {

561  <5><6><7><8>cstt8 <9> cstt9
562  <10><10.95>cstt10
563  <12><14.4><17.28><20.74><24.88>cstt12
564  }{}
565 \DeclareFontShape{IL2}{cmtt}{m}{it}

566 {

567  <5> <6> <7> <8> <9>
568  <10> <10.95> <12> <14.4> <17.28> <20.74> <24.88> csitt10
569  }{}
570 \DeclareFontShape{IL2}{cmtt}{m}{sl}

571 {

572  <5> <6> <7> <8> <9>
573  <10> <10.95> <12> <14.4> <17.28> <20.74> <24.88> cssltt10
574  }{}
575 \DeclareFontShape{IL2}{cmtt}{m}{sc}

576 {

577  <5> <6> <7> <8> <9>

578  <10> <10.95> <12> <14.4> <17.28> <20.74> <24.88> cstcsc10
579  }{}
580 \DeclareFontShape{IL2}{cmtt}{m}{ui}
581 {<->ssub * cmtt/m/it}{}
582 \DeclareFontShape{IL2}{cmtt}{bx}{n}
583 {<->ssub * cmtt/m/n}{}
584 \DeclareFontShape{IL2}{cmtt}{bx}{it}
585 {<->ssub * cmtt/m/it}{}
586 \DeclareFontShape{IL2}{cmtt}{bx}{ui}
587 {<->ssub * cmtt/m/it}{}

588 ⟨/il2cmtt⟩ 589 ⟨∗il2cmvtt⟩

590 \ProvidesFile{il2cmvtt.fd}[1997/02/13 CsLaTeX font definitions]
591 \DeclareFontFamily{IL2}{cmvtt}{\hyphenchar\font45 }
592 \DeclareFontShape{IL2}{cmvtt}{m}{n}

593 {

594  <5><6><7><8><9><10><10.95>
595  <12><14.4><17.28><20.74><24.88>csvtt10
596  }{}
597 \DeclareFontShape{IL2}{cmvtt}{m}{it}

598 {

599  <5><6><7><8><9><10><10.95>
600  <12><14.4><17.28><20.74><24.88>csvtti10
601  }{}

602 ⟨/il2cmvtt⟩ 603 ⟨∗il2lcmss⟩

604 \ProvidesFile{il2lcmss.fd}[1997/02/13 v2.2a CsLaTeX slide font definitions]
605 \DeclareFontFamily{IL2}{lcmss}{\hyphenchar\font45 }
606 \DeclareFontShape{IL2}{lcmss}{m}{n}{%
607 <7><8><10><12><13.82><16.59><19.907><23.89><28.66><34.4><41.28>
608 lcsss8
609  }{}
610 \DeclareFontShape{IL2}{lcmss}{m}{In}{%
611 <7><8><10><12><13.82><16.59><19.907><23.89><28.66><34.4><41.28>
612 ilcsss8
613  }{}
614 \DeclareFontShape{IL2}{lcmss}{m}{sl}{%
615  <13.82><16.59><19.907><23.89><28.66><34.4><41.28>
616 lcsssi8
617  }{}
618 \DeclareFontShape{IL2}{lcmss}{m}{Isl}{%
619  <13.82><16.59><19.907><23.89><28.66><34.4><41.28>
620 ilcsssi8
621  }{}
622 \DeclareFontShape{IL2}{lcmss}{m}{it}{%
623  <-> sub * lcmss/m/sl }{}
624 \DeclareFontShape{IL2}{lcmss}{m}{Iit}{%
625  <-> sub * lcmss/m/Isl }{}
626 \DeclareFontShape{IL2}{lcmss}{bx}{n}{%
627  <13.82><16.59><19.907><23.89><28.66><34.4><41.28>
628 lcsssb8
629  }{}
630 \DeclareFontShape{IL2}{lcmss}{bx}{In}{%
631  <13.82><16.59><19.907><23.89><28.66><34.4><41.28>
632 ilcsssb8
633  }{}
634 \DeclareFontShape{IL2}{lcmss}{m}{ui}{%
635  <-> sub * cmr/m/ui }{}

636 \DeclareFontShape{IL2}{lcmss}{bx}{ui}{%
637  <-> sub * cmr/m/ui }{}

638 ⟨/il2lcmss⟩ 639 ⟨∗il2lcmtt⟩

640 \ProvidesFile{il2lcmtt.fd}[1997/02/13 CsLaTeX slide font definitions]
641 \DeclareFontFamily{IL2}{lcmtt}{\hyphenchar\font\m@ne}
642 \DeclareFontShape{IL2}{lcmtt}{m}{n}{%
643  <13.82><16.59><19.907><23.89><28.66><34.4><41.28>
644 cstt8
645  }{}
646 \DeclareFontShape{IL2}{lcmtt}{m}{In}{%
647  <13.82><16.59><19.907><23.89><28.66><34.4><41.28>
648 icstt8
649  }{}
650 \DeclareFontShape{IL2}{lcmtt}{m}{it}{%
651  <13.82><16.59><19.907><23.89><28.66><34.4><41.28>
652 csitt10
653  }{}
654 \DeclareFontShape{IL2}{lcmtt}{m}{ui}{%
655  <-> sub * cmtt/m/it }{}
656 \DeclareFontShape{IL2}{lcmtt}{bx}{ui}{%
657  <-> sub * cmtt/m/it }{}

658 ⟨/il2lcmtt⟩

5 Czech and slovak style files—czech.sty and slovak.sty

659 ⟨czech⟩\ifx\dateczech\undefined\else\expandafter\endinput\fi
660 ⟨slovak⟩\ifx\dateslovak\undefined\else\expandafter\endinput\fi

661 ⟨∗czech | slovak⟩

662 \newif\iflte
663 \ifx\documentclass\undefined\else\ltetrue\fi
664 \iflte
665 \NeedsTeXFormat{LaTeX2e}

666 ⟨/czech | slovak⟩ 667 ⟨∗czech⟩

668 \ProvidesFile{czech.sty}[1997/04/19 v2.0 CSTeX czech style]
669 \else
670 \message{Document Style Option `czech' ver. 2.0 <April 1997>.}

671 ⟨/czech⟩ 672 ⟨∗slovak⟩

673 \ProvidesFile{slovak.sty}[1997/04/19 v2.0 CSTeX slovak style]
674 \else
675 \message{Document Style Option `slovak' ver. 2.0 <April 1997>.}

676 ⟨/slovak⟩ 677 ⟨∗czech | slovak⟩

678 \edef\origcatcodeat{\the\catcode`\@}\catcode`\@=11
679 \let\providecommand=\def
680 \let\protect=\relax
681 \fi
682 \iflte
683 \DeclareOption{IL2}{\def\encodingdefault{IL2}\let\il@opt=\relax}
684 \DeclareOption {T1}{\def\encodingdefault {T1}}
685 \DeclareOption{OT1}{\def\encodingdefault{OT1}}
686 \DeclareOption{nosplit}{\standardhyphens}
687 \DeclareOption{split}{\splithyphens}
688 \DeclareOption{nocaptions}{\let\cap@unchgd=\relax}
689 \DeclareOption{cstex}{\let\quote@hding=\relax}

690 \ExecuteOptions{IL2}
691 \ProcessOptions
692 \def\dms#1#2{\DeclareMathSymbol{#1}{\mathalpha}{letters}{#2}}

For IL2, encoding of mathematical fonts is changed to IL2 (this encoding is equivalent with OT1 in slots 0–127). This allow using diacritical characters in some font families. Another consequence is that document does not use CM fonts where there are equivalents in CS-fonts, so there is no need to have both bitmaps of CM-font and CS-font.

693 \ifx\il@opt\relax
694  \let\cs@tmp=\@font@warning \let\@font@warning=\@gobble
695  \SetSymbolFont{operators}{normal}{IL2}{cmr}{m}{n}
696  \let\@font@warning=\cs@tmp \let\cs@tmp=\undefined
697  \SetMathAlphabet\mathbf{normal}{IL2}{cmr}{bx}{n}
698  \SetMathAlphabet\mathrm{normal}{IL2}{cmr}{m}{n}
699  \SetMathAlphabet\mathsf{normal}{IL2}{cmss}{m}{n}
700  \SetMathAlphabet\mathtt{normal}{IL2}{cmtt}{m}{n}
701  \dms{^^e1}{"E1}\dms{^^c1}{"C1}\dms{^^e8}{"E8}\dms{^^c8}{"C8}
702  \dms{^^ef}{"EF}\dms{^^cf}{"CF}\dms{^^e9}{"E9}\dms{^^c9}{"C9}
703  \dms{^^ec}{"EC}\dms{^^cc}{"CC}\dms{^^ed}{"ED}\dms{^^cd}{"CD}
704  \dms{^^b5}{"B5}\dms{^^a5}{"A5}\dms{^^f2}{"F2}\dms{^^d2}{"D2}
705  \dms{^^f3}{"F3}\dms{^^d3}{"D3}\dms{^^f8}{"F8}\dms{^^d8}{"D8}
706  \dms{^^b9}{"B9}\dms{^^a9}{"A9}\dms{^^bb}{"BB}\dms{^^ab}{"AB}
707  \dms{^^fa}{"FA}\dms{^^da}{"DA}\dms{^^f9}{"F9}\dms{^^d9}{"D9}
708  \dms{^^fd}{"FD}\dms{^^dd}{"DD}\dms{^^be}{"BE}\dms{^^ae}{"AE}
709 \fi
710 \else
711 \def\gobble#1{}
712 \def\DeclareRobustCommand#1#2{\expandafter\def\csname
713  @\expandafter\gobble\string#1\endcsname{#2}
714  \edef#1{\noexpand\protect\expandafter\noexpand\csname
715  @\expandafter\gobble\string#1\endcsname}}
716 \ifx\ou\undefined \def\ou{\accent23u} \fi
717 \def\cstieon{\def~{\ifmmode\else\penalty\@M\ \fi}}
718 \def\cstieoff{\def~{\penalty\@M \ }}
719 \def\temp#1#2#3:{#1#2}
720 \edef\tempa{\string c\string s}
721 \edef\tempb{\expandafter\temp\fontname\tenrm:}
722 \ifx\tempa\tempb
723  \chardef\clqq=254  \sfcode254=0
724  \chardef\crqq=255  \sfcode255=0
725  \chardef\flqq=158  \sfcode158=0
726  \chardef\frqq=159  \sfcode159=0
727  \def\ogonek #1{\setbox0\hbox{#1}\ifdim\ht0=1ex\accent157 #1%
728  \else{\ooalign{\unhbox0\crcr\hss\char157}}\fi}
729  \chardef\promile=141
730  \def\DeclareTextCommandDefault#1#2{}
731 \else
732  \let\DeclareTextCommandDefault=\DeclareRobustCommand
733 \fi
734 \fi

735 ⟨/czech | slovak⟩

Czech captions

736 ⟨∗czech⟩

737 \def\captionsczech{%
738 \def\abstractname{Abstrakt}%
739 \def\appendixname{P\v{r}\'{\i}loha}%
740 \def\bibname{Literatura}%
741 \def\ccname{Na v\v{e}dom\'{\i}:}

742 \def\chaptername{Kapitola}%
743 \def\contentsname{Obsah}%
744 \def\enclname{P\v{r}\'{\i}loha}%
745 \def\figurename{Obr\'azek}%
746 \def\headpagename{Strana}%
747 \def\headtoname{Komu}%
748 \def\indexname{Rejst\v{r}\'{\i}k}%
749 \def\listfigurename{Seznam obr\'azk\r{u}}%
750 \def\listtablename{Seznam tabulek}%
751 \def\partname{\v{C}\'ast}%
752 \def\prefacename{P\v{r}edmluva}%
753 \def\proofname{D\r{u}kaz}%
754 \def\seename{viz}%
755 \def\alsoseename{viz tak\'e}%
756 \def\refname{Reference}%
757 \def\tablename{Tabulka}}

758 ⟨/czech⟩

Slovak captions

759 ⟨∗slovak⟩

760 \def\captionsslovak{%
761 \def\prefacename{Predhovor}%
762 \def\refname{Literat\'ura}%
763 \def\appendixname{Dodatok}%
764 \def\listfigurename{Zoznam obr\'azkov}%
765 \def\listtablename {Zoznam tabuliek}%
766 \def\indexname{Register}%
767 \def\tablename{Tabu\v{l}ka}%
768 \def\partname{\v{C}as\v{t}}%
769 \def\enclname{Pr\'{\i}loha}%
770 \def\headtoname{Pre}%
771 \def\alsoname{vi\v{d} tie\v{z}}%
772 \def\abstractname{Abstrakt}%
773 \let\bibname\refname
774 \def\chaptername{Kapitola}%
775 \def\contentsname{Obsah}%
776 \def\figurename{Obr.}%
777 \def\ccname{cc.}%
778 \def\pagename{Str.}%
779 \def\seename{vi\v{d}}}

780 ⟨/slovak⟩

English captions

781 ⟨∗czech | slovak⟩

782 \providecommand\captionsenglish{%
783 \def\prefacename{Preface}%
784 \def\refname{References}%
785 \def\abstractname{Abstract}%
786 \def\bibname{Bibliography}%
787 \def\chaptername{Chapter}%
788 \def\appendixname{Appendix}%
789 \def\contentsname{Contents}%
790 \def\listfigurename{List of Figures}%
791 \def\listtablename{List of Tables}%
792 \def\indexname{Index}%
793 \def\figurename{Figure}%
794 \def\tablename{Table}%
795 \def\partname{Part}%
796 \def\enclname{encl}%
797 \def\ccname{cc}%

798 \def\headtoname{To}%
799 \def\pagename{Page}%
800 \def\headpagename{Page}%
801 \def\prefacename{Preface}%
802 \def\seename{see}%
803 \def\alsoname{see also}}

804 ⟨/czech | slovak⟩

Czech date

805 ⟨∗czech⟩

806 \def\dateczech{%
807 \def\today{\number\day. \ifcase\month\or ledna\or \'unora\or
808  b\v{r}ezna\or dubna\or kv\v{e}tna\or \v{c}ervna\or \v{c}ervence\or
809  srpna\or z\'a\v{r}\'\i\or \v{r}\'{\i}jna\or listopadu\or
810  prosince\fi \space\number\year}}

Switching on the czech captions and date (if not requested else)

811 \iflte\ifx\cap@unchgd\undefined\captionsczech\dateczech\fi\fi

812 ⟨/czech⟩

Slovak date

813 ⟨∗slovak⟩

814 \def\dateslovak{\def\today{\number\day.~\ifcase\month\or
815 janu\'ara\or febru\'ara\or marca\or apr\'{\i}la\or m\'aja\or j\'una\or
816 j\'ula\or augusta\or septembra\or okt\'obra\or
817 novembra\or decembra\fi
818 \space\number\year}}

Switching on the slovak captions and date (if not requested else)

819 \iflte\ifx\cap@unchgd\undefined\captionsslovak\dateslovak\fi\fi

820 ⟨/slovak⟩

English (19th April 1997) and US (April 19, 1997) date

821 ⟨∗czech | slovak⟩

822 \providecommand\dateUSenglish{\def\today{\ifcase\month\or
823  January\or February\or March\or April\or May\or June\or July\or
824  August\or September\or October\or November\or December\fi
825  \space\number\day, \number\year}}
826 \providecommand\dateenglish{\def\today{\ifcase\day\or 1st\or 2nd\or
827  3rd\or 4th\or 5th\or 6th\or 7th\or 8th\or 9th\or 10th\or 11th\or
828  12th\or 13th\or 14th\or 15th\or 16th\or 17th\or 18th\or 19th\or
829  20th\or 21st\or 22nd\or 23rd\or 24th\or 25th\or 26th\or 27th\or
830  28th\or 29th\or 30th\or 31st\fi ~\ifcase\month\or January\or
831  February\or March\or April\or May\or June\or July\or August\or
832  September\or October\or November\or December\fi \space
833  \number\year}}

7bin cs quotes (for OT1)

834 \DeclareTextCommandDefault\clqq{\ifhmode
835 \edef\@SF{\spacefactor\the\spacefactor}\else\let\@SF\empty\fi\leavevmode
836 \set@low@box{''}\box0
837 \setbox1=\hbox{l\/}\dimen1=\wd1
838 \setbox1=\hbox{l}\advance\dimen1 by -\wd1
839 \ifdim\dimen1>0pt \kern.01em\else\kern-.04em\fi
840 \nobreak\hskip0pt\@SF\relax}
841 \DeclareTextCommandDefault\crqq{\edef\@SF{\spacefactor\the\spacefactor}%
842 \kern-.07em\hbox{``}\kern.07em\@SF\relax}

Lowers the box to ... of ...

843 \def\set@low@box#1{\setbox2=\hbox{,}\setbox0=\hbox{#1}%
844 \dimen0=\ht0 \advance\dimen0 by -\ht2
845 \setbox0=\hbox{\lower\dimen0 \box0}\ht0=\ht2\dp0=\dp2}

Single cs quotes

846 \DeclareRobustCommand\clq{\leavevmode\set@low@box{\char'047 }%
847 \box0 \kern.04em\nobreak\hskip0pt\relax}
848 \DeclareRobustCommand\crq{{\edef\@SF{\spacefactor\the\spacefactor}%
849 \nobreak\hskip0pt\char'140 \kern.17em\@SF\relax}}

Command for making quotes

850 \DeclareRobustCommand\uv{\bgroup\aftergroup\closequotes\leavevmode\clqq
851 \let\next=}
852 \def\closequotes{\unskip\crqq\relax}

Active quotes (“” -¿ ,,“)

853 {\catcode`\'=\active
854 \gdef\csprimeson{\catcode96=\active \catcode39=\active
855 \def\pr@m@s{\ifx'\next\let\nxt\pr@@@s \else\ifx^\next\let\nxt\pr@@@t
856  \else\let\nxt\egroup\fi\fi \nxt}}}
857 \def\csprimesoff{\catcode96=12 \catcode39=12
858 \def\pr@m@s{\ifx'\next\let\nxt\pr@@@s \else\ifx^\next\let\nxt\pr@@@t
859  \else\let\nxt\egroup\fi\fi \nxt}}
860 {\csprimeson
861 \gdef'{\ifmmode\let\n@xt=\mathprim@\else\let\n@xt=\textprim@\fi\n@xt}%

We need to preserve function of macro for converting f′′′ -¿ f′′′ in math mode.

862 \gdef\mathprim@{^\bgroup\prim@s}%
863 \gdef\textprim@{\futurelet\nxt\rightprim@}%
864 \gdef\rightprim@{\ifx'\nxt\let\next=\douvr@\else\let\next\crq\fi\next}%
865 \gdef`{\futurelet\nxt\leftprim@}%
866 \gdef\leftprim@{\ifx`\nxt\let\next=\douvl@\else\let\next=\clq\fi\next}}
867 \def\douvr@{\crqq\let\next= }%
868 \def\douvl@{\clqq\let\next= }%

Hyphenation and language switching

869 \iflte

870 ⟨/czech | slovak⟩

871 ⟨czech⟩\AtBeginDocument{\czech}
872 ⟨slovak⟩\AtBeginDocument{\slovak}

873 ⟨∗czech | slovak⟩

874 \frenchspacing
875 \DeclareTextSymbol{\clqq}{T1}{"12}
876 \DeclareTextCommand{\crqq}{T1}{{\edef\@SF{\spacefactor
877  \the\spacefactor}\char16\@SF\relax}}
878 \else
879 \def\setthisl@nguage#1#2#3#4{%
880 \ifx#2\undefined \immediate\write\sixt@@n{Warning: I do not speak #1,
881  (the style is not inputted)}\else
882  \ifx#4\undefined \immediate\write\sixt@@n{Warning: I do not speak #1,
883  (the hyphenation patterns are not included)}\else
884  #2#3#4\fi\fi}
885 \def\selectlanguage#1{\language #1\relax
886  \ifcase #1\relax           \dateUSenglish \captionsenglish \ehyph \or
887  \setthisl@nguage{german}   \dategerman    \captionsgerman  \ghyph \or
888  \setthisl@nguage{austrian} \dateaustrian  \captionsgerman  \ahyph \or
889  \setthisl@nguage{french}   \datefrench    \captionsfrench  \fhyph \or
890  \setthisl@nguage{english}  \dateenglish   \captionsenglish \ehyph \or
891  \setthisl@nguage{czech}    \dateczech     \captionsczech   \chyph \or
892  \setthisl@nguage{slovak}   \dateslovak    \captionsslovak  \shyph \fi}
893 \def\originalTeX{\selectlanguage{\USenglish}\csname cmaccents\endcsname}
894 \def\czechTeX{\selectlanguage{\czech}\csname csaccents\endcsname}
895 \def\slovakTeX{\selectlanguage{\slovak}\csname csaccents\endcsname}
896 \catcode`\@=\origcatcodeat

897 ⟨czech⟩\czechTeX 898 ⟨slovak⟩\slovakTeX

899 \fi
900 \endinput

901 ⟨/czech | slovak⟩